# Optimal estimation for regression discontinuity design

### v1

@@ Appendix B Minimax estimation for the average treatment effect
In addition, we have $b_{+}(\bm{w}_{+},\tilde{\bm{\theta}}_{+}(\theta_{0,+}))\geq b_{+}(\tilde{\bm{w}}_{+},\tilde{\bm{\theta}}_{+}(\theta_{0,+}))\geq 0$. Hence, we have
$\mathrm{MSE}_{ate}\left(\bm{w}_{+},\bm{w}_{-},\tilde{\bm{\theta}}_{+}(\theta_{0,+}),\tilde{\bm{\theta}}_{-}(\theta_{0,-})\right)\ \geq\ \mathrm{MSE}_{ate}\left(\tilde{\bm{w}}_{+},\bm{w}_{-},\tilde{\bm{\theta}}_{+}(\theta_{0,+}),\tilde{\bm{\theta}}_{-}(\theta_{0,-})\right)$
Hence, if $w_{j,+}<w_{j+1,+}$, then we can reduce the maximum MSE by exchanging $w_{j,+}$ for $w_{j+1,+}$. Similar argument holds for $\bm{w}_{-}$. Therefore, for any $\bm{w}_{+}\in\mathcal{W}_{+}$ and $\bm{w}_{-}\in\mathcal{W}_{-}$, there exists $(\tilde{\bm{w}}_{+},\tilde{\bm{w}}_{-})\in\mathcal{W}_{+}\times\mathcal{W}_{-}$ such that $\tilde{w}_{1,+}\geq\cdots\geq\tilde{w}_{n_{+},+}$, $\tilde{w}_{1,-}\geq\cdots\geq\tilde{w}_{n_{-},-}$, and
@@
$=$ $2\theta_{j,+}\left\{\left(\sum_{i\neq j}w_{i,+}\theta_{i,+}-\theta_{0,+}\right)-b_{-}(\bm{w}_{-},\bm{\theta}_{-})\right\}+w_{j,+}/2.$
@@
Because $g(\bm{w};\theta_{0,+},\theta_{0,-})$ is convex with respect to $\bm{w}$ for all $\theta_{0,+}$ and $\theta_{0,-}$, $\overline{g}(\bm{w})$ is also convex. Therefore, we can solve the minimax problem by minimizing $\overline{g}(\bm{w})$ subject to $\bm{w}\in\mathcal{W}_{+}^{1}\times\mathcal{W}_{-}^{1}$.
@@ Appendix C Confidence intervals with general bounded outcomes
C.1 One-sided confidence intervals
@@
We consider one-sided $100\cdot(1-\alpha)\%$ CIs $\hat{\tau}-\gamma,\infty)$ satisfying
$\inf_{\bm{p}\in\mathcal{P},Q\in\mathcal{Q}(\bm{p})}P_{\bm{p},Q}\left(\tau_{0}\in[\hat{\tau}-\gamma,\infty)\right)\ \geq\ 1-\alpha,\ \text{ or }\ \sup_{\bm{p}\in\mathcal{P},Q\in\mathcal{Q}(\bm{p})}P_{\bm{p},Q}\left(\hat{\tau}-\tau_{0}>\gamma\right)\ \leq\ \alpha,$
@@
We construct a CI by using an upper bound on $\sup_{\bm{p}\in\mathcal{P},Q\in\mathcal{Q}(\bm{p})}P_{\bm{p},Q}\left(\hat{\tau}-\tau_{0}>\gamma\right)$. Define
${\rm Bias}_{\bm{p}}(\hat{\tau})\equiv E_{\bm{p}}[\hat{\tau}]-\tau_{0}=\sum_{i=1}^{n_{+}}w_{i,+}\left(p_{i,+}-\frac{1}{2}\right)-\sum_{i=1}^{n_{-}}w_{i,-}\left(p_{i,-}-\frac{1}{2}\right)-\tau_{0}$
@@
$P_{\bm{p},Q}\left(\hat{\tau}-\tau_{0}>\gamma\right)$
@@
$\sup_{\bm{p}\in\mathcal{P},Q\in\mathcal{Q}(\bm{p})}P_{\bm{p},Q}\left(\hat{\tau}-\tau_{0}>\gamma\right)\ \leq\ \exp\left(-\frac{2(\gamma-\overline{{\rm Bias}}_{\mathcal{P}}(\hat{\tau}))^{2}}{\sum_{i=1}^{n_{+}}w_{i,+}^{2}+\sum_{i=1}^{n_{-}}w_{i,-}^{2}}\right).$
@@
C.2 Two-sided confidence intervals
We consider two-sided $100\cdot(1-\alpha)\%$ CIs $\hat{\tau}-\gamma,\hat{\tau}+\gamma]$ satisfying
$\inf_{\bm{p}\in\mathcal{P},Q\in\mathcal{Q}(\bm{p})}P_{\bm{p},Q}\left(\tau_{0}\in[\hat{\tau}-\gamma,\hat{\tau}+\gamma]\right)\ \geq\ 1-\alpha,\ \text{ or }\ \sup_{\bm{p}\in\mathcal{P},Q\in\mathcal{Q}(\bm{p})}P_{\bm{p},Q}\left(|\hat{\tau}-\tau_{0}|>\gamma\right)\ \leq\ \alpha.$
@@ C.2 Two-sided confidence intervals
$\sup_{\bm{p}\in\mathcal{P},Q\in\mathcal{Q}(\bm{p})}P_{\bm{p},Q}\left(|\hat{\tau}-\tau_{0}|>\gamma\right)\ \leq\ 2\exp\left(-\frac{2(\gamma-\overline{{\rm Bias}}_{\mathcal{P}}(\hat{\tau}))^{2}}{\sum_{i=1}^{n_{+}}w_{i,+}^{2}+\sum_{i=1}^{n_{-}}w_{i,-}^{2}}\right).$
@@
$P_{\bm{p},Q}\left(|\hat{\tau}-\tau_{0}|>\gamma\right)$
@@
$\sup_{\bm{p}\in\mathcal{P},Q\in\mathcal{Q}(\bm{p})}P_{\bm{p},Q}\left(|\hat{\tau}-\tau_{0}|>\gamma\right)$
@@
Note that $\overline{\pi}(\gamma)$ is tighter than a naive upper bound:
$\overline{\pi}(\gamma)\ <\ 2\exp\left(-\frac{2(\gamma-\overline{{\rm Bias}}_{\mathcal{P}}(\hat{\tau}))^{2}}{\sum_{i=1}^{n_{+}}w_{i,+}^{2}+\sum_{i=1}^{n_{-}}w_{i,-}^{2}}\right).$
This implies that $\gamma^{*}<\overline{{\rm Bias}}_{\mathcal{P}}(\hat{\tau})+\left(\frac{\log(1/(\alpha/2))\left(\sum_{i=1}^{n_{+}}w_{i,+}^{2}+\sum_{i=1}^{n_{-}}w_{i,-}^{2}\right)}{2}\right)^{1/2}$.
@@ Appendix D Optimal Weights in Gaussian Models
We use Donoho (1994)’s results to derive optimal weights that solve the minimax problem (3.1) and show that the weights satisfy $\sum_{i=1}^{n}w_{i}=1$ and $w_{i}\geq 0$ for all $i$. See Armstrong and Kolesár (2021) for an application of Donoho (1994)’s results in a related setting. Our Gaussian setting falls into the framework of Donoho (1994). Specifically, in the notation of Donoho (1994), we observe $\mathbf{y}$ of the form $\mathbf{y}=K\mathbf{x}+\mathbf{z}$ with $\mathbf{x}\in\mathbf{X}$. Here, $\mathbf{y}=(Y_{1}/\sigma_{1},\ldots,Y_{n}/\sigma_{n})^{\prime}$, $\mathbf{z}\sim N(0,I_{n})$, where $I_{n}$ is an $n\times n$ identity matrix, $\mathbf{x}=\bm{\theta}$, $\mathbf{X}=\Theta_{g}$, and $K\mathbf{x}=(\theta_{1}/\sigma_{1},\ldots,\theta_{n}/\sigma_{n})^{\prime}$. The parameter of interest is the linear functional $L\mathbf{x}=\theta_{0}$. We derive an affine estimator that minimizes the maximum MSE among all affine estimators (i.e, estimators of form $\hat{\theta}_{0}=c+\bm{w}^{\prime}\mathbf{y}=c+\sum_{i=1}^{n}(w_{i}/\sigma_{i})Y_{i}$ with $c\in\mathbb{R}$ and $\bm{w}\in\mathbb{R}^{n}$).
To specialize the results in Donoho (1994) to our setting, define the modulus of continuity of $L$:
@@
The results in Donoho (1994) (in particular, the arguments in the proof of Theorem 1) then yield the following result. Let $\varepsilon_{0}>0$ be a solution to $\frac{(\varepsilon/2)^{2}}{(\varepsilon/2)^{2}+1}=\frac{\varepsilon\omega^{\prime}(\varepsilon)}{\omega(\epsilon)}$ and let $\bm{\theta}_{\varepsilon_{0}}$ solve (D.1) at $\varepsilon=\varepsilon_{0}$. Then, the following estimator minimizes the maximum MSE among all affine estimators:
@@
To prove (iii), we apply Lemma D.1 in Supplemental Appendix D of Armstrong and Kolesár (2018). Our setting falls into their framework where $f=\bm{\theta}$, $\mathcal{F}=\mathcal{G}=\Theta_{g}$, $Kf=(\theta_{1}/\sigma_{1},\ldots,\theta_{n}/\sigma_{n})^{\prime}$, and $Lf=\theta_{0}$ in their notation. To apply their Lemma D.1, let $\iota\in\mathbb{R}^{n+1}$ denote the vector of ones. Then, we have $\iota\in\Theta_{g}$, $L\iota=1$, and $\bm{\theta}_{\varepsilon}+c\iota\in\Theta_{g}$ for all $c\in\mathbb{R}$. By their Lemma D.1, $\omega(\cdot)$ is differentiable at $\varepsilon>0$ with

### v2

@@ Appendix B Minimax estimation for the average treatment effect
In addition, noting that
$b_{+}(\bm{w}_{+},\bm{\theta}_{+})-b_{+}(\tilde{\bm{w}}_{+},\bm{\theta}_{+})=(w_{j,+}-w_{j+1,+})(\theta_{j,+}-\theta_{j+1,+})$
and
$b_{+}(\tilde{\bm{w}}_{+},\bm{\theta}_{+})=\sum_{i=1}^{n_{+}}\tilde{w}_{i,+}(\theta_{i,+}-\theta_{0,+})-\left(1-\sum_{i=1}^{n_{+}}\tilde{w}_{i,+}\right)\theta_{0,+},$
we have $b_{+}(\bm{w}_{+},\tilde{\bm{\theta}}_{+}(\theta_{0,+}))\geq b_{+}(\tilde{\bm{w}}_{+},\tilde{\bm{\theta}}_{+}(\theta_{0,+}))\geq 0$ for any $\theta_{0,+}\in[-1/2,0]$. Also, $b_{-}(\bm{w}_{-},\tilde{\bm{\theta}}_{-}(\theta_{0,-}))\leq 0$ for any $\theta_{0,-}\in[0,1/2]$. Hence, we have
$\mathrm{MSE}_{ate}\left(\bm{w}_{+},\bm{w}_{-},\tilde{\bm{\theta}}_{+}(\theta_{0,+}),\tilde{\bm{\theta}}_{-}(\theta_{0,-})\right)\ \geq\ \mathrm{MSE}_{ate}\left(\tilde{\bm{w}}_{+},\bm{w}_{-},\tilde{\bm{\theta}}_{+}(\theta_{0,+}),\tilde{\bm{\theta}}_{-}(\theta_{0,-})\right).$
Therefore, if $w_{j,+}<w_{j+1,+}$, then we can reduce the maximum MSE by exchanging $w_{j,+}$ for $w_{j+1,+}$. Similar argument holds for $\bm{w}_{-}$. Therefore, for any $\bm{w}_{+}\in\mathcal{W}_{+}$ and $\bm{w}_{-}\in\mathcal{W}_{-}$, there exists $(\tilde{\bm{w}}_{+},\tilde{\bm{w}}_{-})\in\mathcal{W}_{+}\times\mathcal{W}_{-}$ such that $\tilde{w}_{1,+}\geq\cdots\geq\tilde{w}_{n_{+},+}$, $\tilde{w}_{1,-}\geq\cdots\geq\tilde{w}_{n_{-},-}$, and
@@
$=$ $2\theta_{j,+}\left\{\left(\sum_{i\neq j}w_{i,+}\theta_{i,+}-\theta_{0,+}\right)-b_{-}(\bm{w}_{-},\bm{\theta}_{-})\right\}+w_{j,+}/2$
$=$ $2\theta_{j,+}\left\{\left(\sum_{i\neq j}w_{i,+}(\theta_{i,+}-\theta_{0,+})-\left(1-\sum_{i\neq j}w_{i,+}\right)\theta_{0,+}\right)-b_{-}(\bm{w}_{-},\bm{\theta}_{-})\right\}+w_{j,+}/2.$
@@
Because $g(\bm{w};\theta_{0,+},\theta_{0,-})$ is convex with respect to $\bm{w}$ for all $\theta_{0,+}$ and $\theta_{0,-}$, $\overline{g}(\bm{w})$ is also convex. We can solve the minimax problem by minimizing $\overline{g}(\bm{w})$ subject to $\bm{w}\in\mathcal{W}_{+}^{1}\times\mathcal{W}_{-}^{1}$.
@@ Appendix C Confidence intervals with general bounded outcomes
C.1 One-sided confidence interval
@@
We consider a one-sided $100\cdot(1-\alpha)\%$ CI $\hat{\tau}-\gamma,\infty)$ satisfying
$\inf_{\bm{p}\in\mathcal{P},Q\in\mathcal{Q}(\bm{p})}P_{\bm{p},Q}\left(\tau\in[\hat{\tau}-\gamma,\infty)\right)\ \geq\ 1-\alpha,\ \text{ or }\ \sup_{\bm{p}\in\mathcal{P},Q\in\mathcal{Q}(\bm{p})}P_{\bm{p},Q}\left(\hat{\tau}-\tau>\gamma\right)\ \leq\ \alpha,$
@@
We construct a CI using an upper bound on $\sup_{\bm{p}\in\mathcal{P},Q\in\mathcal{Q}(\bm{p})}P_{\bm{p},Q}\left(\hat{\tau}-\tau>\gamma\right)$. Define
${\rm Bias}_{\bm{p}}(\hat{\tau})\equiv E_{\bm{p}}[\hat{\tau}]-\tau=\sum_{i=1}^{n_{+}}w_{i,+}\left(p_{i,+}-\frac{1}{2}\right)-\sum_{i=1}^{n_{-}}w_{i,-}\left(p_{i,-}-\frac{1}{2}\right)-\tau$
@@
$P_{\bm{p},Q}\left(\hat{\tau}-\tau>\gamma\right)$
@@
$\sup_{\bm{p}\in\mathcal{P},Q\in\mathcal{Q}(\bm{p})}P_{\bm{p},Q}\left(\hat{\tau}-\tau>\gamma\right)\ \leq\ \exp\left(-\frac{2(\gamma-\overline{{\rm Bias}}_{\mathcal{P}}(\hat{\tau}))^{2}}{\sum_{i=1}^{n_{+}}w_{i,+}^{2}+\sum_{i=1}^{n_{-}}w_{i,-}^{2}}\right).$
@@
$\tilde{\bm{p}}_{+}(0)$ $\equiv$ $\left(0,\min\{C\|R_{1,+}\|,1\},\ldots,\min\{C\|R_{n_{+},+}\|,1\}\right)^{\prime},$
$\tilde{\bm{p}}_{-}(1)$ $\equiv$ $\left(1,\max\{1-C\|R_{1,-}\|,0\},\ldots,\max\{1-C\|R_{n_{-},-}\|,0\}\right)^{\prime}.$
Therefore, $\gamma^{*}$ has a closed-form expression.
C.2 Two-sided confidence interval
We consider a two-sided $100\cdot(1-\alpha)\%$ CI $\hat{\tau}-\gamma,\hat{\tau}+\gamma]$ satisfying
$\inf_{\bm{p}\in\mathcal{P},Q\in\mathcal{Q}(\bm{p})}P_{\bm{p},Q}\left(\tau\in[\hat{\tau}-\gamma,\hat{\tau}+\gamma]\right)\ \geq\ 1-\alpha,\ \text{ or }\ \sup_{\bm{p}\in\mathcal{P},Q\in\mathcal{Q}(\bm{p})}P_{\bm{p},Q}\left(|\hat{\tau}-\tau|>\gamma\right)\ \leq\ \alpha.$
[…]
$\sup_{\bm{p}\in\mathcal{P},Q\in\mathcal{Q}(\bm{p})}P_{\bm{p},Q}\left(|\hat{\tau}-\tau|>\gamma\right)\ \leq\ 2\exp\left(-\frac{2(\gamma-\overline{{\rm Bias}}_{\mathcal{P}}(\hat{\tau}))^{2}}{\sum_{i=1}^{n_{+}}w_{i,+}^{2}+\sum_{i=1}^{n_{-}}w_{i,-}^{2}}\right).$ (C.1)
@@ C.2 Two-sided confidence interval
$P_{\bm{p},Q}\left(|\hat{\tau}-\tau|>\gamma\right)$
@@
$\sup_{\bm{p}\in\mathcal{P},Q\in\mathcal{Q}(\bm{p})}P_{\bm{p},Q}\left(|\hat{\tau}-\tau|>\gamma\right)$
@@
Note that $\overline{\pi}(\gamma)$ is a tighter upper bound than the bound given in (C.1), since
$\overline{\pi}(\gamma)$ $\leq\max_{b\in[0,\overline{{\rm Bias}}_{\mathcal{P}}(\hat{\tau})]}\exp\left(-\frac{2(\gamma-b)^{2}}{\sum_{i=1}^{n_{+}}w_{i,+}^{2}+\sum_{i=1}^{n_{-}}w_{i,-}^{2}}\right)$
$\quad\quad+\max_{b\in[0,\overline{{\rm Bias}}_{\mathcal{P}}(\hat{\tau})]}\exp\left(-\frac{2(\gamma+b)^{2}}{\sum_{i=1}^{n_{+}}w_{i,+}^{2}+\sum_{i=1}^{n_{-}}w_{i,-}^{2}}\right)$
$=\exp\left(-\frac{2(\gamma-\overline{{\rm Bias}}_{\mathcal{P}}(\hat{\tau}))^{2}}{\sum_{i=1}^{n_{+}}w_{i,+}^{2}+\sum_{i=1}^{n_{-}}w_{i,-}^{2}}\right)+\exp\left(-\frac{2\gamma^{2}}{\sum_{i=1}^{n_{+}}w_{i,+}^{2}+\sum_{i=1}^{n_{-}}w_{i,-}^{2}}\right)$
$\leq 2\exp\left(-\frac{2(\gamma-\overline{{\rm Bias}}_{\mathcal{P}}(\hat{\tau}))^{2}}{\sum_{i=1}^{n_{+}}w_{i,+}^{2}+\sum_{i=1}^{n_{-}}w_{i,-}^{2}}\right).$
This implies that $\gamma^{*}\leq\overline{{\rm Bias}}_{\mathcal{P}}(\hat{\tau})+\left(\frac{\log(1/(\alpha/2))\left(\sum_{i=1}^{n_{+}}w_{i,+}^{2}+\sum_{i=1}^{n_{-}}w_{i,-}^{2}\right)}{2}\right)^{1/2}$, which leads to a less conservative CI.
@@ Appendix D Optimal Weights in Gaussian Models
We use donoho1994’s results to derive optimal weights that solve the minimax problem (3.1) and show that the weights satisfy $\sum_{i=1}^{n}w_{i}=1$ and $w_{i}\geq 0$ for all $i$. See Armstrong2021ATE for an application of donoho1994’s results in a related setting. Our Gaussian setting falls into the framework of donoho1994. Specifically, in the notation of donoho1994, we observe $\mathbf{y}$ of the form $\mathbf{y}=K\mathbf{x}+\mathbf{z}$ with $\mathbf{x}\in\mathbf{X}$. Here, $\mathbf{y}=(Y_{1}/\sigma_{1},\ldots,Y_{n}/\sigma_{n})^{\prime}$, $\mathbf{z}\sim N(0,I_{n})$, where $I_{n}$ is an $n\times n$ identity matrix, $\mathbf{x}=\bm{\theta}$, $\mathbf{X}=\Theta_{g}$, and $K\mathbf{x}=(\theta_{1}/\sigma_{1},\ldots,\theta_{n}/\sigma_{n})^{\prime}$. The parameter of interest is the linear functional $L\mathbf{x}=\theta_{0}$. We derive an affine estimator that minimizes the maximum MSE among all affine estimators (i.e, estimators of form $\hat{\theta}_{0}=c+\bm{w}^{\prime}\mathbf{y}=c+\sum_{i=1}^{n}(w_{i}/\sigma_{i})Y_{i}$ with $c\in\mathbb{R}$ and $\bm{w}\in\mathbb{R}^{n}$).
To specialize the results in donoho1994 to our setting, define the modulus of continuity of $L$:
@@
The results in donoho1994 (in particular, the arguments in the proof of Theorem 1) then yield the following result. Let $\varepsilon_{0}>0$ be a solution to $\frac{(\varepsilon/2)^{2}}{(\varepsilon/2)^{2}+1}=\frac{\varepsilon\omega^{\prime}(\varepsilon)}{\omega(\epsilon)}$ and let $\bm{\theta}_{\varepsilon_{0}}$ solve (D.1) at $\varepsilon=\varepsilon_{0}$. Then, the following estimator minimizes the maximum MSE among all affine estimators:
@@
To prove (iii), we apply Lemma D.1 in Supplemental Appendix D of Armstrong.Kolesar2018. Our setting falls into their framework where $f=\bm{\theta}$, $\mathcal{F}=\mathcal{G}=\Theta_{g}$, $Kf=(\theta_{1}/\sigma_{1},\ldots,\theta_{n}/\sigma_{n})^{\prime}$, and $Lf=\theta_{0}$ in their notation. To apply their Lemma D.1, let $\iota\in\mathbb{R}^{n+1}$ denote the vector of ones. Then, we have $\iota\in\Theta_{g}$, $L\iota=1$, and $\bm{\theta}_{\varepsilon}+c\iota\in\Theta_{g}$ for all $c\in\mathbb{R}$. By their Lemma D.1, $\omega(\cdot)$ is differentiable at $\varepsilon>0$ with